Current of 10 ampere is passed through solution of CuSO4 for 32 min 10 sec.If 896 ml of Oxygen is liberated at STP then what will be efficiency of the cell.

Grade:12

1 Answers

Vikas TU
7 years ago
Dear Candidate,
I already answered this qstn.
plz check this below:
 
from faraday’s law,
moles x valence factor = i*t/F
moles = i*t/F*n
          = 10*1930/2*96485
         = 0.100 approx.
Moles given at STP = 0.896/22.4 = 0.04
efficiency = (0.1-0.04)/0.1  * 100 => 60%